OpenStudy (anonymous):

The moon is a sphere with radius of 567 km. Determine an equation for the ellipse if the distance of the satellite from the surface of the moon varies from 900 km to 169 km.

OpenStudy (anonymous):

x^2/900 + y^2/169 = 1 x^2/1467^2 + y^2/736^2= 1 x^2/736^2 + y^2/1467 = 1 x^2/169 + y^2/900 = 1

OpenStudy (anonymous):

huh?

OpenStudy (shamim):

the equation is x^2/a^2+y^2/b^2=1

OpenStudy (anonymous):

i know but those were my answer choices

OpenStudy (shamim):

a=567+900 b=567+169

OpenStudy (anonymous):

ohh thanks

OpenStudy (anonymous):

i just realized how easy that was

OpenStudy (shamim):

wc